OpenStudy (anonymous):

Find an expression for the arc length of the upper half of a circle of radius 25 on the interval [0,t] where (0

OpenStudy (anonymous):

@satellite73 yo can u try this

OpenStudy (anonymous):

finally !!!

OpenStudy (anonymous):

i wrote 25 pi/2 but it was wrong then i wrote t*pi/2 that was wrong too

OpenStudy (anonymous):

we have to use arclength formula just incase

OpenStudy (anonymous):

i forget is it someting like \[\int_a^b\sqrt{1+(f'(x))^2}\]

myininaya (myininaya):

ya

OpenStudy (anonymous):

ya thats correct

myininaya (myininaya):

25 is a huge radius

OpenStudy (anonymous):

in this case the circle with radius 25 has the equation \[f(x)=\sqrt{25^2-t^2}\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

it is rather large, but not as large as i am old

OpenStudy (anonymous):

hey i dont judge a question by its size @myininaya

OpenStudy (anonymous):

=P

OpenStudy (anonymous):

take the derivative, get \[\frac{t}{\sqrt{25^2-t^2}}\]

myininaya (myininaya):

and you could actually compare this answer by doing it the easy algebraic way

myininaya (myininaya):

you know to check the answer

OpenStudy (anonymous):

that is what i started to do actually but it required too much thinking

OpenStudy (anonymous):

^

OpenStudy (anonymous):

square it, add one, and take the square root of the result halas

myininaya (myininaya):

C=2pi*r C=2pi*25=50pi 0 to 25 means we are looking at a quarter of the whole circle C/4=50pi/4 50pi/4=25pi/2 hmmm...I got 25pi/2 like you did

myininaya (myininaya):

is it 0<t<25 or -25<t<25?

OpenStudy (anonymous):

0 to 25

myininaya (myininaya):

oh it says find the expression and doesn't say to evaluate

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

lol and its confusing because interval is from 0 to t and 0<t<25...so...

myininaya (myininaya):

I think we she do the integral for 0 to t

myininaya (myininaya):

I think that inequality is just saying where t has to be less than 25

OpenStudy (anonymous):

basically we have to find such a t that it the arc length was from 0 to t (t<25) that we can get the answer

myininaya (myininaya):

\[\int\limits_{0}^{t} \sqrt{1+(\frac{-t}{25^2-t^2})^2 } dx\] and satellite already found the derivative of f

myininaya (myininaya):

oops those t's are x's

myininaya (myininaya):

except the limit t

OpenStudy (anonymous):

the question just asks for the expression

myininaya (myininaya):

\[\int\limits\limits_{0}^{t} \sqrt{1+(\frac{-x}{25^2-x^2})^2 } dx \]

OpenStudy (anonymous):

still a typo there dear

myininaya (myininaya):

whatever I missed the square root

OpenStudy (anonymous):

\[\int\limits\limits_{0}^{t} \sqrt{1+(\frac{-x}{\sqrt{25^2-x^2}})^2 } dx\]

myininaya (myininaya):

\[\int\limits\limits\limits_{0}^{t} \sqrt{1+(\frac{-x}{\sqrt{25^2-x^2}})^2 } dx \]

myininaya (myininaya):

too many latex symbols to type

OpenStudy (anonymous):

or \[\sqrt{1+\frac{x^2}{25^2-x^2}}\]

OpenStudy (anonymous):

alright you guys i got it right

OpenStudy (anonymous):

i had it right before just that i was evaluating from 0 to 25 when i was just spposed to sub in for t...sigh didnt read...